OpenStudy (anonymous):

Use division to determine whether the ratios form a proportion. 12 : 44 and 3 : 11 A. No. They do not form a proportion. B. Yes. They form a proportion.

OpenStudy (texaschic101):

12/44 = 3/11 .272727 = .272727 yup....thats a proportion

OpenStudy (texaschic101):

think about it...12/44 reduces to 3/11...and proportions are nothing but equivalent fractions

OpenStudy (triciaal):

all ratios are proportions anyway, check if 12/44 can be reduced factors of 12 4*3 factors of 44 4*11 common factor 4 can be reduced 12/44 = 3/11 so yes B

OpenStudy (anonymous):

Thanks :)